OpenStudy (anonymous):

Solve for x. -ax + 3b > 5 x > the quantity 3 times b minus 5 all over negative a x > the quantity 5 minus 3 times b all over negative a x < the quantity 3 times b plus 5 all over negative a x < the quantity negative 3 times b plus 5 all over negative a

OpenStudy (anonymous):

you cannot solve this

OpenStudy (anonymous):

\[-ax+3b>5\] \[-ax>5-3b\] divide by \(-1\) get \[ax<3b-5\] but you cannot go further

TheSmartOne (thesmartone):

\[x<\frac{ 3b-5 }{ a }\] Yes you can @satellite73

OpenStudy (anonymous):

the reason you cannot divide both sides by \(a\) is because \(a\) is a variable, and could be either positive or negative if it is positive you leave the inequality alone if it is negative, you have to switch the inequality where does this question come from?

OpenStudy (anonymous):

@TheSmartOne like hell you cannot divide an inequality by a variable

TheSmartOne (thesmartone):

Ok.

OpenStudy (anonymous):

@luckyjeans please tell me where this question comes from? an idiot math teacher, or an idiot on line system, like FLVS

TheSmartOne (thesmartone):

But 5-3b=-3b+5 right??

OpenStudy (anonymous):

oops yes, you are right, they are the same still want to know the origin of this question

TheSmartOne (thesmartone):

oh and then you divided by -1 nevermind...

TheSmartOne (thesmartone):

same here...

OpenStudy (anonymous):

\[\huge \begin{array}l\color{red}{\normalsize\text{w}}\color{orange}{\normalsize\text{h}}\color{#9c9a2e}{\normalsize\text{a}}\color{green}{\normalsize\text{t}}\color{blue}{\normalsize\text{ }}\color{purple}{\normalsize\text{i}}\color{purple}{\normalsize\text{s}}\color{red}{\normalsize\text{ }}\color{orange}{\normalsize\text{w}}\color{#9c9a2e}{\normalsize\text{r}}\color{green}{\normalsize\text{o}}\color{blue}{\normalsize\text{n}}\color{purple}{\normalsize\text{g}}\color{purple}{\normalsize\text{ }}\color{red}{\normalsize\text{w}}\color{orange}{\normalsize\text{i}}\color{#9c9a2e}{\normalsize\text{t}}\color{green}{\normalsize\text{h}}\color{blue}{\normalsize\text{ }}\color{purple}{\normalsize\text{t}}\color{purple}{\normalsize\text{h}}\color{red}{\normalsize\text{i}}\color{orange}{\normalsize\text{s}}\color{#9c9a2e}{\normalsize\text{ }}\color{green}{\normalsize\text{q}}\color{blue}{\normalsize\text{u}}\color{purple}{\normalsize\text{e}}\color{purple}{\normalsize\text{s}}\color{red}{\normalsize\text{t}}\color{orange}{\normalsize\text{i}}\color{#9c9a2e}{\normalsize\text{o}}\color{green}{\normalsize\text{n}}\color{blue}{\normalsize\text{?}}\color{purple}{\normalsize\text{}}\end{array}\]